9) Twice the tens digit of a number increased by the units digit is 22. If the digits are reversed, the new number is 45 less than the original.
10) Three times the units digit is 2 less than the tens digit. If the digits are reversed, the new number is 54 less than the original.
(THANK YOU)

Answer by mananth(16949)

x tens digit: y units digit
2x+y=22..........1
10y+x= 10x+y-45
9y-9x=-45
y-x=-5...........2
multiply equation (2) by 2
we get 2y-2x==-10
add this equation to (1)
3y=12
y=4
...
2x+y=22
plug y
2x+4=22
2x=18
x=9
number is 94